# Primes and absolutely or non-absolutely irreducible elements in atomic domains

### v1

@@ 6.1. Absolutely Irreducible Elements in Krull monoids
We note in passing that the statements of the previous lemma are further equivalent to $\mathcal{B}(G)$ being half-factorial.
@@
For the converse, suppose that every irreducible of $H$ is absolutely irreducible. By Theorem 2, every irreducible of $\mathcal{B}(G)$ must be absolutely irreducible. Thus $\left|G\right|\leq 2$ by Lemma 6.5. To show that $G$ is trivial it therefore suffices to show $G\not\cong\mathbb{Z}/2\mathbb{Z}$.
Suppose that $G=\{0,g\}\cong\mathbb{Z}/2\mathbb{Z}$. By (b) of [15, Theorem 2.5.4.1] there must exist at least two distinct non-zero divisorial prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of non-trivial class $[\mathfrak{p}]=[\mathfrak{q}]=g$. Therefore $\mathfrak{p}^{2}$, $\mathfrak{q}^{2}$ and $\mathfrak{p\cdot_{v}q}$ are all principal and generated by irreducibles $u$, $v$, $w$, say, $\mathfrak{p}^{2}=uH$, $\mathfrak{q}^{2}=vH$ and $\mathfrak{pq}=wH$. Now $w^{2}\sim uv$ shows that $w$ is not absolutely irreducible. ∎
@@
For the converse, suppose that every absolutely irreducible element is prime. Suppose that $H$ is not factorial. Then there exist non-prime irreducibles. Let $u\in H$ be such an irreducible with $\operatorname{supp}(uH)$ minimal among all irreducibles. Then $u$ is absolutely irreducible by Proposition 6.4, and therefore prime, a contradiction. ∎
@@ 7. Rings with irreducible elements that are all absolutely irreducible, but none of them prime
It follows from [17, Theorem 8] that there exists a Dedekind domain $D$ with class group $G=\mathbb{Z}^{n}$ and $G_{0}$ precisely the set of classes containing prime ideals. There exists a transfer homomorphism $\varphi\colon D\setminus\{0\}\to\mathcal{B}(G_{0})$ (see Theorem 1). In particular, the domain $D$ is not half-factorial and an element $a\in D$ is irreducible if and only if the corresponding zero-sum sequence $\varphi(a)\in\mathcal{B}(G_{0})$ is irreducible.
Since $0\notin G_{0}$, the trivial ideal class of $D$ contains no prime ideals. Hence $D$ contains no prime element. Finally, every irreducible element of $\mathcal{B}(G_{0})$ is square-free, and so Theorem 2 implies that every irreducible element of $D$ is absolutely irreducible. ∎
@@ 8. Rings with all irreducible elements absolutely irreducible but not all prime
There exists a Dedekind domain $D$ that is not half-factorial, contains a prime element, and such that all of its irreducibles elements are absolutely irreducible.

### v2

@@ 7.1. Consequences for Krull monoids
The statements of the previous lemma are further equivalent to $\mathcal{B}(G)$ being half-factorial [15, Theorem 3.4.11.5].
@@
For the converse, suppose that every irreducible of $H$ is absolutely irreducible. By Theorem 7.2, every irreducible of $\mathcal{B}(G)$ must be absolutely irreducible. Thus $\left|G\right|\leq 2$ by Corollary 7.3. To show that $G$ is trivial it therefore suffices to show $G\not\cong\mathbb{Z}/2\mathbb{Z}$.
Suppose that $G=\{0,g\}\cong\mathbb{Z}/2\mathbb{Z}$. By (b) of [15, Theorem 2.5.4.1] there must exist at least two distinct non-zero divisorial prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of non-trivial class $\mathfrak{p}]=[\mathfrak{q}]=g$. Therefore $\mathfrak{p}^{2}$, $\mathfrak{q}^{2}$ and $\mathfrak{p\cdot_{v}q}$ are all principal and generated by irreducibles $u$, $v$, $w$, say, $\mathfrak{p}^{2}=uH$, $\mathfrak{q}^{2}=vH$ and $\mathfrak{p\cdot_{v}q}=wH$. Now $w^{2}\sim uv$ shows that $w$ is not absolutely irreducible. ∎
@@
For the converse, suppose that every absolutely irreducible element is prime. Suppose that $H$ is not factorial. Then there exist non-prime irreducibles. Let $u\in H$ be a non-prime irreducible with $\operatorname{supp}(uH)$ minimal among all non-prime irreducibles. Then $\operatorname{supp}(uH)$ is in fact minimal among all irreducibles: if $p\in H$ is a prime, then $pH$ is a divisorial prime ideal and hence $\operatorname{supp}(pH)=\{pH\}\subseteq\operatorname{supp}(uH)$ would imply that $pH$ appears in the factorization of $uH$ into divisorial prime ideals, meaning $p\mid u$, which is impossible since $u$ is a non-prime irreducible. Then $u$ is absolutely irreducible by Proposition 7.1, and therefore prime, a contradiction. ∎
@@ 8. Rings with irreducible elements that are all absolutely irreducible, but none of them prime
It follows from [17, Theorem 8] that there exists a Dedekind domain $D$ with class group $G=\mathbb{Z}^{n}$ and $G_{0}$ precisely the set of classes containing prime ideals. There exists a transfer homomorphism $\varphi\colon D\setminus\{0\}\to\mathcal{B}(G_{0})$ (see Theorem 2.9). In particular, the domain $D$ is not half-factorial and an element $a\in D$ is irreducible if and only if the corresponding zero-sum sequence $\varphi(a)\in\mathcal{B}(G_{0})$ is irreducible.
Since $0\notin G_{0}$, the trivial ideal class of $D$ contains no prime ideals. Hence $D$ contains no prime element. Finally, every irreducible element of $\mathcal{B}(G_{0})$ is square-free, and so Theorem 7.2 implies that every irreducible element of $D$ is absolutely irreducible. ∎
@@ 9. Rings with all irreducible elements absolutely irreducible but not all prime
There exists a Dedekind domain $D$ that is not half-factorial, contains a prime element, and such that all of its irreducible elements are absolutely irreducible.
@@
Acknowledgment. The authors wish to thank the second of two anonymous referees for contributing an additional example of a ring containing primes and both absolutely and non-absolutely irreducible elements to Section 6, namely, the one in Example 6.1.